# Stanford CA 94305, USA

### v2

@@ 5 Randomized quasi-Monte Carlo
Two important issues are how the error in the CLT (6) decreases with $n$ and how quickly the coverage of the standard student $t$ confidence interval (9) approaches $1-\alpha$ as $n\to\infty$. These are most often studied through moment quantities. Let $\bar{Y}$ be the average of $n$ IID random variables $Y_{i}$ that have mean $\mu$, variance $\sigma^{2}>0$ and a finite third moment. Then by the Berry-Esseen theorem, there exists a constant $C<\infty$ such that
@@
$\gamma=\frac{\mathbb{E}((Y-\mu)^{3})}{\sigma^{3}}\quad\text{and}\quad\kappa=\frac{\mathbb{E}((Y-\mu)^{4})}{\sigma^{3}}-3.$
These are known as the skewness and kurtosis of $Y$. A Gaussian random variable has $\gamma=\kappa=0$. If $\kappa<\infty$ and $Y_{i}$ are IID random variables with the same distribution as $Y$ then $\hat{\mu}=\bar{Y}=(1/n)\sum_{i=1}^{n}Y_{i}$ has skewness $\gamma/\sqrt{n}$ and kurtosis $\kappa/n$ [61]. Then the skewness of $\bar{Y}$ approaches 00 more slowly than the kurtosis does. That is consistent with the third moment but not the fourth appearing in the Berry-Esseen bound.
@@
This error has been well studied by researchers in the 1980s, especially Peter Hall [38, 39], in the context of bootstrap confidence intervals described below. A very interesting finding is that for IID sampling of $n$ observations, the coverage error in the standard interval is commonly $O(1/n)$ which is better than the RMSE of $O(n^{-1/2})$ for $\hat{\mu}$. The UQ is then more accurate than the estimate whose uncertainty it quantifies. The one-sided coverage errors $\Pr\bigl{(}\mu\leqslant\bar{Y}-{s}{t^{1-\alpha/2}_{(n-1)}/\sqrt{n}}\bigr{)}-\alpha/2$ and $\Pr\bigl{(}\mu\geqslant\bar{Y}+{s}t^{1-\alpha/2}_{(n-1)}/{\sqrt{n}}\bigr{)}-\alpha/2$ both converge at the $O(n^{-1/2})$ rate but a fortunate cancellation yields an $O(n^{-1})$ rate for (15).
@@
Of the 7 continuous distributions in [65], the bootstrap $t$ got close to nominal coverage in 6 of them. The exception was the lognormal distribution, where the bootstrap $t$ got only about 90% coverage by $n=20$, which was noticeably higher than all the other 8 methods apart from an alternative formulation of the bootstrap $t$.
@@
Hall [39] gives some asymptotic expansions for the coverage error in confidence intervals for the mean. His Table 1 includes the percentile bootstrap, the bootstrap $t$ and the normal theory interval (8) that uses Gaussian quantiles instead of $t$ quantiles. Surprisingly, he does not include the standard Student $t$ intervals. The coverage errors for two sided confidence intervals are
@@
Normal theory: $\qquad(1/n)\varphi(z^{1-\alpha/2})\bigl{[}\phantom{-}0.14\kappa-2.12\gamma^{2}-3.35\bigr{]}+O(1/{n^{2}}),$
Percentile: $\qquad(1/n)\varphi(z^{1-\alpha/2})\bigl{[}-0.72\kappa-0.37\gamma^{2}-3.35\bigr{]}+O(1/{n^{2}}),\quad\text{and}$
@@
Bootstrap $t$: $\qquad(1/n)\varphi(z^{1-\alpha/2})\bigl{[}-2.84\kappa+4.25\gamma^{2}\,\,\,\phantom{-3.35}\bigr{]}+O(1/{n^{2}}),$
@@
where $\varphi$ is the probability density function of the standard Gaussian distribution. Hall’s names for these methods are, ‘Norm’, ‘BACK’ and ‘STUD’ respectively. His assumptions include $8$ finite moments and a distribution for $Y_{i}$ whose support is not contained within an arithmetic sequence. His table is for nominal coverage $1-2\alpha$ whereas we ordinarily target coverage $1-\alpha$, but this only affects the scaling of the lead term, and not the qualitative consequences of skewness and kurtosis. Hall’s formulas for $n=18$ were very close to the average coverage for $16\leqslant n\leqslant 20$ in [65, Table 6], except for the lognormal distribution.
@@
Hall’s formulas show an advantage for the bootstrap t. The term $4.25\gamma^{2}$ is positive, so it works to increase coverage. It also has no intercept, while the other two methods have an intercept of $-3.35$ which works to decrease their coverage. Simulations of coverage levels usually show that nonparametric confidence intervals have a coverage level that appoaches the desired one from below as $n\to\infty$. That is, undercoverage is more common than overcoverage. While Hall does not give a formula for the standard student based interval we know that it is exact for Gaussian data, and Gaussian data have $\gamma=\kappa=0$. We thus expect the corresponding formula to have a zero intercept, though not necessarily with the same coefficients for skewness and kurtosis that the normal theory interval has.
@@
A large value of the kurtosis is known to give the $t$ statistic a more concentrated distribution (closer to zero) than the nominal $t_{(n-1)}$ distribution that it has for Gaussian data. That makes it less likely to exceed a threshold like $t^{1-\alpha/2}_{(n-1)}$ which in turn makes the confidence interval more likely to include $\mu$. An explanation and survey of this work, going back nearly 100 years, appears in [14]. What happens is that very large values of $Y_{i}$ inflate the denominator $s/\sqrt{n}$ more than the numerator $\bar{Y}-\mu$. In the extreme, if we send $Y_{1}\to\pm\infty$ then $t\to\pm 1$. The consequence is that for distributions with large kurtosis, the usual 95% confidence intervals, based on $|t|\leqslant t_{(n-1)}^{0.975}$ can cover the mean more often than the nominal level, because $t_{(n-1)}^{0.975}\geqslant 1.96$.
@@
Now for the RQMC context, we replace $n$ by $R$ and $Y_{i}$ by $\hat{\mu}_{n}^{(r)}$. Only a little is known about the skewness and kurtosis of $\hat{\mu}_{n}$ for various RQMC methods and much of that knowledge is recent. Let these be $\gamma_{n}$ and $\kappa_{n}$, respectively.
@@
First, there is an old result [57] proving a CLT as $n\to\infty$ for the scrambled nets from [66] when the underlying digital net has $t=0$, under smoothness conditions on $f$. That setting then has $\gamma_{n}\to 0$ and $\kappa_{n}\to 0$ and we expect the standard intervals should then have low coverage error. The nets of Faure [31] have $t=0$ but the more widely used ones based on sequences from Sobol’ [81] only have $t=0$ for $d\leqslant 2$ [64, Table 4.1]. On the other hand, randomly shifted lattice rules are known to produce $\hat{\mu}$ that does not follow a CLT [58]. The distribution of $\hat{\mu}$ has instead a density represented by a spline curve.
@@
The simulations in [52] investigated three different confidence interval methods for RQMC: the standard interval, the percentile bootstrap and the bootstrap $t$. The number of replicates was $R\in\{5,10,20,30\}$ in keeping with the desire to keep $R$ small. There were 5 different RQMC methods. Two of them were lattice rules from LatNet Builder [51] using random shifts modulo one, with and without the baker transformation of [42]. The other three RQMC methods were Sobol’s digital nets [81] using the direction numbers from [47]. They were randomized with either a digital shift, a matrix scramble of [59] plus digital shift or a nested uniform scramble from [66]. The sample sizes were $n=2^{m}$ for $m\in\{6,8,10,12,14\}$. There were $6$ integrands with known $\mu$ chosen to mix cases that are easy and difficult for RQMC, varying in their levels of smoothness and varying in the extent to which they have low effective dimension. See [52] for a discussion. An integrand that is easy to integrate well is not necessarily one that makes the confidence interval problem easy. All of those integrands were constructed to allow varying dimension $d$. The simulation used $d\in\{4,8,16,32\}$.
@@
Of the 2400 cases, the percentile bootstrap was found to fail 1689 times. This is not surprising given the results in [65]. The bootstrap t was found to fail 81 times. The standard interval failed only 3 times. None of those failures were for $R=10$. Figure 2 of [52] erroneosly shows one such failure. That happened because the plotting command to set up the figure’s axes mistakenly did not use type =“n” in the call to the plot function in R.
@@
Based on the results in [52], the current best recommendation for 95% RQMC confidence intervals is to use $R\geqslant 10$ independent replications along with the standard student $t$ based confidence interval, along with one’s preferred RQMC method. Next we relate this finding to the understanding of confidence intervals from Hall’s formulas.
@@
From the data in [52] it was possible to compute sample values of the skewness and kurtosis in each distribution of $\hat{\mu}$ based on $10{,}000$ evaluations for each integrand, dimension $d$, sample size $n$ and RQMC method. Inspection of those estimated values revealed that most of the RQMC distributions had modest sample skewnesses and many of them had very large sample kurtoses. A large kurtosis presents a difficulty for the bootstrap $t$ and an advantage for the standard interval as discussed above. The modest skewness removes a disadvantage for the standard interval and it removes an advantage for the bootstrap $t$. Figure 2 of [52] shows that in some instances with extremely large kurtosis, the standard confidence interval had coverage well above the nominal 95% level. This is to be expected based on [14].
@@
The empirical findings motivated subsequent work in [74] which shows that for random linear scrambling of a digital net in base $2$ with a digital shift, the skewness of $\hat{\mu}_{n}$ is $\gamma_{n}=O(n^{\epsilon})$ for any $\epsilon>0$, so it is almost $O(1)$. Furthermore, under a model with randomly chosen generator matrices the skewness is $O(n^{-1/2+\epsilon})$. The kurtosis of $\hat{\mu}$ from random linear scrambling of a digital net in base $2$ is known to diverge to $\infty$ as $n\to\infty$ for an analytic function on $[0,1]$ by a finding in [73, Section 3]. The reasoning is as follows. There is an event of probability $\Omega(1/n)$ that gives a squared error of $\Omega(1/n^{2})$ (for smooth $f^{\prime}$ that does not integrate to zero). That same event on its own contributes $\Omega(n^{-5})$ to the fourth power of the error. The expected squared error is $O(n^{-3+\epsilon})$ and so $\kappa_{n}+3=\Omega(n^{-5})/O(n^{-6+2\epsilon})=\Omega(n^{1-2\epsilon})$. An integrand on $[0,1]^{d}$ for $d>1$ with a smooth contribution from a one dimensional main effect will have the same diverging kurtosis. Having $\kappa_{n}\to\infty$ completely rules out a CLT for $\hat{\mu}_{n}$ from matrix scrambling with a digital shift.
The empirical findings of modest skewness and potentially very large kurtosis have not been established for the other RQMC methods. The other methods in [52] did not show large skewness empirically. Many of the histograms of $\hat{\mu}$ had non-Gaussian but nearly symmetric distributions. There is an elegant orthant symmetry argument in [27] showing conservativeness of the standard $t$ based confidence interval for symmetrically distributed data with or without high kurtosis. Note that symmetry implies zero skewness (given a finite third absolute moment) but zero skewness does not imply symmetry.
@@
The bias-corrected accelerated bootstrap (BCa) of [28] was not included in the simulations of either [52] or [65]. Hall [39] prefers the bootstrap $t$, while DiCiccio and Efron [19] and others prefer the BCa. Table 1 of [39] includes an entry ‘ABC’ for a method that is very close to BCa. It has a coverage expression of $-2.63\kappa+3.11\gamma^{2}-3.35$. The constant $-3.35$ and the negative coefficient for $\kappa$ suggest that BCa is not well suited to the RQMC estimates that can have very large kurtosis.

### v3

@@ 5 Randomized quasi-Monte Carlo
Two important issues are how the error in the CLT (6) decreases with $n$ and how quickly the coverage of the standard Student’s $t$ confidence interval (9) approaches $1-\alpha$ as $n\to\infty$. These are most often studied through moment quantities. Let $\bar{Y}$ be the average of $n$ IID random variables $Y_{i}$ that have mean $\mu$, variance $\sigma^{2}>0$ and a finite third moment. Then by the Berry-Esseen theorem, there exists a constant $C<\infty$ such that
@@
$\gamma=\frac{\mathbb{E}((Y-\mu)^{3})}{\sigma^{3}}\quad\text{and}\quad\kappa=\frac{\mathbb{E}((Y-\mu)^{4})}{\sigma^{4}}-3.$
These are known as the skewness and kurtosis of $Y$. A Gaussian random variable has $\gamma=\kappa=0$. If $\kappa<\infty$ and $Y_{i}$ are IID random variables with the same distribution as $Y$ then $\hat{\mu}=\bar{Y}=(1/n)\sum_{i=1}^{n}Y_{i}$ has skewness $\gamma/\sqrt{n}$ and kurtosis $\kappa/n$ [71]. Then the skewness of $\bar{Y}$ approaches 00 more slowly than the kurtosis does. A rapidly vanishing kurtosis is consistent with the third moment appearing in the Berry-Esseen bound but the fourth moment not appearing there.
@@
This error has been well studied by researchers in the 1980s, especially Peter Hall [44, 45], in the context of bootstrap confidence intervals described below. A very interesting finding is that for IID sampling of $n$ observations, the coverage error in the standard interval is commonly $O(1/n)$ which is better than the RMSE of $O(n^{-1/2})$ for $\hat{\mu}$. The UQ achieves a better convergence rate than the estimate whose uncertainty it quantifies. The one-sided coverage errors $\Pr\bigl{(}\mu\leqslant\bar{Y}-{s}{t^{1-\alpha/2}_{(n-1)}/\sqrt{n}}\bigr{)}-\alpha/2$ and $\Pr\bigl{(}\mu\geqslant\bar{Y}+{s}t^{1-\alpha/2}_{(n-1)}/{\sqrt{n}}\bigr{)}-\alpha/2$ both converge at the $O(n^{-1/2})$ rate but a fortunate cancellation yields an $O(n^{-1})$ rate for (15).
@@
Of the 7 continuous distributions in [75], the bootstrap $t$ got close to nominal coverage in 6 of them. The exception was the lognormal distribution, where the bootstrap $t$ got only about 90% coverage by $n=20$. Even that was noticeably higher than all the other 8 methods apart from an alternative formulation of the bootstrap $t$.
@@
Hall [45] gives some asymptotic expansions for the coverage error in confidence intervals for the mean. His Table 1 includes the percentile bootstrap, the bootstrap $t$ and the normal theory interval (8) that uses Gaussian quantiles instead of $t$ quantiles. Surprisingly, he does not include the standard Student’s $t$ intervals; for those see [80]. The coverage errors for two sided confidence intervals are
@@
Normal theory: $\qquad(2/n)\varphi(z^{1-\alpha/2})\bigl{[}\phantom{-}0.14\kappa-2.12\gamma^{2}-3.35\bigr{]}+O(1/{n^{2}}),$
@@
Student’s $t$: $\qquad(2/n)\varphi(z^{1-\alpha/2})\bigl{[}\phantom{-}0.14\kappa-2.12\gamma^{2}\,\,\,\phantom{-3.42}\bigr{]}+O(1/{n^{2}}),$
Percentile: $\qquad(2/n)\varphi(z^{1-\alpha/2})\bigl{[}-0.72\kappa-0.37\gamma^{2}-3.35\bigr{]}+O(1/{n^{2}}),\quad\text{and}$
Bootstrap $t$: $\qquad(2/n)\varphi(z^{1-\alpha/2})\bigl{[}-2.84\kappa+4.25\gamma^{2}\,\,\,\phantom{-3.42}\bigr{]}+O(1/{n^{2}}),$
where $\varphi$ is the probability density function of the standard Gaussian distribution. Hall’s names for the normal theory, percentile and bootstrap $t$ methods are, ‘Norm’, ‘BACK’ and ‘STUD’ respectively.
@@
Hall’s assumptions include $8$ finite moments and a distribution for $Y_{i}$ whose support is not contained within an arithmetic sequence. His table is for nominal coverage $1-2\alpha$ whereas we ordinarily target coverage $1-\alpha$, but this only affects the scaling of the lead term, and not the qualitative consequences of skewness and kurtosis. Hall’s formulas for $n=18$ were very close to the average coverage for $16\leqslant n\leqslant 20$ in [75, Table 6], except for the lognormal distribution.
@@
The coverage error formulas show an advantage for the bootstrap $t$. The term $4.25\gamma^{2}$ is positive, so it works to increase coverage. It also has no intercept, while the normal theory and percentile methods have a negative intercept which works to decrease their coverage. Simulations of coverage levels usually show that nonparametric confidence intervals have a coverage level that approaches the desired one from below as $n\to\infty$. That is, undercoverage is more common than overcoverage. Student’s $t$ intervals have a positive coefficient for kurtosis, a negative coefficient for squared skewness and no intercept.
@@
The distribution of the $t$ statistic under non-normality is much studied. For IID data from a distribution symmetric about $\mu$ with heavier tails than the Gaussian, there is a tendency for confidence intervals based on the $t$ statistic to be conservative, though precise statements of this phenomenon require some extra steps. An explanation and survey of this work, going back nearly 100 years, appears in [17]. At a high level, higher kurtosis in $Y_{i}$ brings a lower kurtosis in $t$, so that extreme thresholds are exceeded by $|t|$ less often. What happens is that very large values of $|Y_{i}-\mu|$ inflate the denominator $s/\sqrt{n}$ more than the numerator $\bar{Y}-\mu$. In the extreme, if we send $Y_{1}\to\pm\infty$ then $t\to\pm 1$. The consequence is that for distributions with large kurtosis, the usual 95% confidence intervals, based on $|t|\leqslant t_{(n-1)}^{0.975}$ can cover the mean more often than the nominal level, because $t_{(n-1)}^{0.975}\geqslant 1.96$.
To switch from folklore to precise statements requires some additional care and caveats. Here are a few of those results. One subtlety is that the tendency to overcoverage holds at customary confidence levels, but not necessarily at lower levels, such as those below 50%. The distribution of $t$ does not depend on $\mu$ so it is studied with $\mu=0$. Then the findings are mostly studied through $\tilde{t}=\tilde{t}_{n}=\sum_{i=1}^{n}Y_{i}/(\sum_{i=1}^{n}Y_{i}^{2})^{1/2}$, with $t=\tilde{t}(n-1)^{1/2}/(n-\tilde{t}^{2})^{1/2}$, a monotone transformation. For even $\nu\geqslant 2$, ${\mathbb{E}}(\tilde{t}^{\nu})\leqslant{\mathbb{E}}(Z^{\nu})$ for $Z\sim\mathcal{N}(0,1)$; see Corollary 1 of [32]. Corollary 2 shows that $\tilde{t}$ has negative kurtosis under conditions that include IID symmetrically distributed $Y_{i}$. If the random variables are distributed as a scale mixture of mean zero Gaussians (i.e., $\mathcal{N}(0,\sigma^{2})$ for random $\sigma$) then the $t$ test is conservative at thresholds above $1.8$ [7]. More precisely: the bound was established for $2\leqslant n\leqslant 18$, with a critical threshold that never went above $1.8$ and decreased with increasing $n>6$. Most of the results are for symmetrically distributed $Y_{i}$, but it is known that the skewness of $t$ is $-2\gamma/\sqrt{n}+O(n^{-3/2})$ where $\gamma$ is the skewness of $Y_{i}$ [71, Equation (1.15)]. For a survey of related results, see [90].
@@
Now we switch to the RQMC context. We replace $n$ by our number $R$ of replicates and $Y_{i}$ by $\hat{\mu}_{n}^{(r)}$. Only a little is known about the skewness and kurtosis of $\hat{\mu}_{n}$ for various RQMC methods and much of that knowledge is recent. Let these be $\gamma_{n}$ and $\kappa_{n}$, respectively.
@@
First, there is an old result [67] proving a CLT as $n\to\infty$ for the scrambled $(t,m,d)$-nets using the scrambling from [76] when the underlying digital net has $t=0$, under smoothness conditions on $f$. That setting then has $\gamma_{n}\to 0$ and $\kappa_{n}\to 0$ and we expect the standard intervals should then have low coverage error. The nets of Faure [36] have $t=0$ but the more widely used ones based on sequences from Sobol’ [94] only have $t=0$ for $d\leqslant 2$ [74, Table 4.1]. On the other hand, randomly shifted lattice rules are known to produce $\hat{\mu}$ that does not follow a CLT [60]. There the distribution of $\hat{\mu}$ has a density represented by a spline curve.
@@
The simulations in [61] investigated three different confidence interval methods for RQMC: the standard interval, the percentile bootstrap and the bootstrap $t$. The number of replicates was $R\in\{5,10,20,30\}$ in keeping with the desire to keep $R$ small. There were 5 different RQMC methods. Two of them were lattice rules from LatNet Builder [59] using random shifts modulo one, with and without the baker transformation of [48]. The other three RQMC methods used Sobol’s digital nets [94] using the direction numbers from [55]. They were randomized with either a digital shift (see [58]), a matrix scramble of [69] plus digital shift, or the nested uniform scramble from [76]. The sample sizes were $n=2^{m}$ for $m\in\{6,8,10,12,14\}$. There were $6$ integrands with known $\mu$ chosen to mix cases that are easy and difficult for RQMC, varying in their levels of smoothness and varying in the extent to which they have low effective dimension. See [61] for a discussion. An integrand that is easy to integrate well is not necessarily one that makes the confidence interval problem easy. All of those integrands were constructed to allow varying dimension $d$. The simulation used $d\in\{4,8,16,32\}$.
@@
Of the 2400 cases, the percentile bootstrap was found to fail 1689 times. This is not surprising given the results in [75]. The bootstrap $t$ was found to fail 81 times. The standard interval failed only 3 times. None of those failures were for $R=10$. Figure 2 of [61] erroneosly shows one such failure. That happened because the plotting command to set up the figure’s axes mistakenly did not use type =“n” in the call to the plot function in the R language.
@@
Based on the results in [61], the current best recommendation for 95% RQMC confidence intervals is to use $R\geqslant 10$ independent replications along with the standard Student’s $t$ based confidence interval, along with one’s preferred RQMC method. Next we relate this finding to the understanding of confidence intervals from Hall’s formulas.
@@
From the data in [61] it was possible to compute sample values of the skewness and kurtosis in each distribution of $\hat{\mu}$ based on $10{,}000$ evaluations for each integrand, dimension $d$, sample size $n$ and RQMC method. Inspection of those estimated values revealed that most of the RQMC distributions had modest sample skewnesses and many of them had very large sample kurtoses. A large kurtosis presents a difficulty for the bootstrap $t$ and an advantage for the standard interval as discussed above. The modest skewness removes a disadvantage for the standard interval and it removes an advantage for the bootstrap $t$. Figure 2 of [61] shows that in some instances with extremely large kurtosis, the standard confidence interval had coverage well above the nominal 95% level. This consequence of high kurtosis is well known [17] as mentioned above.
@@
The empirical findings motivated subsequent work in [85] which shows that for random linear scrambling of a digital net in base $2$ with a random digital shift, the skewness of $\hat{\mu}_{n}$ is $\gamma_{n}=O(n^{\epsilon})$ for any $\epsilon>0$, so it is almost $O(1)$. Furthermore, under a model with randomly chosen generator matrices the skewness is $O(n^{-1/2+\epsilon})$. The kurtosis of $\hat{\mu}$ from random linear scrambling of a digital net in base $2$ is known to diverge to $\infty$ as $n\to\infty$ for an analytic function on $[0,1]$ by a finding in [84, Section 3]. The reasoning is as follows. There is an event of probability $\Omega(1/n)$ that gives a squared error of $\Omega(1/n^{2})$ (for smooth $f^{\prime}$ that does not integrate to zero). That same event on its own contributes $\Omega(n^{-5})$ to the fourth power of the error. The expected squared error is $O(n^{-3+\epsilon})$ and so $\kappa_{n}+3=\Omega(n^{-5})/O(n^{-6+2\epsilon})=\Omega(n^{1-2\epsilon})$. An integrand on $[0,1]^{d}$ for $d>1$ with a smooth contribution from a one dimensional main effect will have the same diverging kurtosis and the critical event has probability $\Omega(d/n^{2})$. Having $\kappa_{n}\to\infty$ completely rules out a CLT for $\hat{\mu}_{n}$ from matrix scrambling with a digital shift.
The empirical findings of modest skewness and potentially very large kurtosis have not been established for the other RQMC methods. The other methods in [61] did not appear to have large skewness. Only three of 600 cases had $|\hat{\gamma}|>4$. Those had large kurtoses and since the variance of $\hat{\gamma}$ involves sixth moments, they might just be sampling fluctuations. Many of the histograms of $\hat{\mu}$ had non-Gaussian but nearly symmetric distributions.
@@
The bias-corrected accelerated bootstrap (BCa) of [33] was not included in the simulations of either [61] or [75]. Hall [45] prefers the bootstrap $t$, while DiCiccio and Efron [23] prefer the BCa. Table 1 of [80] includes an entry ‘ABC’ for a method that is very close to BCa. The BCa has a coverage expression of $-2.68\kappa+3.17\gamma^{2}-3.42$ (after corrections). The constant $-3.42$ and the $-2.68\kappa$ term both suggest that BCa is not well suited to the RQMC estimates that can have very large kurtosis and minimal skewness.